OpenStudy (anonymous):

The formula for the area of a trapezoid is A=1/2h(b1+b2), where A is the area, h is the height, and b1 and b2 are the bases of the trapezoid. Solve the formula for h .

OpenStudy (texaschic101):

A = 1/2h(b1 + b2) --- multiply both sides by 2 2A = h(b1 + b2) --- divide both sides by (b1 + b2) (2A) / (b1 + b2) = h